OpenStudy (anonymous):

What is the magnitude and direction of GH if G(3, -2) and H(6, 4)?

OpenStudy (calculusfunctions):

If given the endpoints \[A(x _{1},y _{1})andB(x _{2},y _{2})\]Then the magnitude\[|AB|=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]and the direction of vector AB is\[\tan \theta =\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]where θ is the angle between the vector and the horizontal.

OpenStudy (souvik):